# Examining Caregiving Roles to Differentiate the Effects of Using a Mobile App for Community Oversight for Privacy and Security

### v1

@@ 1. Introduction
A recent Pew Research study (Vogels and Anderson, 2019) reported that most adults in the U.S. have significant knowledge gaps regarding their digital privacy and security. The report suggests a lack of understanding of how third party entities access personal information online (Auxier et al., 2019), making users susceptible to privacy and security breaches. This begs the question of how we might close this knowledge gap? Previous studies have explored solutions involving one party overseeing another, such as organizations using Mobile Device Management systems to secure employee data (Hayes et al., 2020) and parents monitoring teens through mobile online safety apps (Wisniewski et al., 2017). These studies suggested that collaborative approaches, rather than unilateral controls, can offer better privacy protection by fostering collective learning and discussion  (Akter et al., 2022). Crowdsourcing has also been proposed as a method to improve individual’s mobile privacy decisions such as managing app permissions to safeguard personal information (Lin et al., 2012; Ismail et al., 2017; Zhang and Xu, 2016; Liu et al., 2019), though it often lacks reliability (Rashidi et al., 2018). An approach that has shown promise is to leverage the social influence of informal networks to infuse expertise and exert influence on privacy and security decisions (Mendel and Toch, 2017). For example, several studies have shown that people tend to trust and follow privacy advice from their trusted circles, often turning to friends and family for guidance on digital privacy and security topics (Mendel and Toch, 2017; Das et al., 2014a, b).
@@
Therefore, networked privacy researchers have called for community based approaches to co-manage digital privacy and security, where individuals in a trusted community can help one another equally make individual and collective privacy and security decisions (Wan et al., 2020; Chouhan et al., 2019; Aljallad et al., 2019). While this body of research confirms that people often rely on their social connections to inform their digital privacy and security decisions, it has yet to explore how these dynamics shift when relationships involve knowledge imbalances or power hierarchies, such as caring responsibilities for children or older adults. Prior studies (Akter et al., 2022, 2023b; Chouhan et al., 2019; Kropczynski et al., 2021b) have shown how communities often consist of hierarchical (i.e. uneven) levels of expertise and knowledge. Individuals with more expertise often take the role of “caregivers” to provide support and guidance to others who are less tech-savvy (Van Parys, 2019; Correa et al., 2015; Kiesler et al., 2000; Poole et al., 2009). Other individuals in the community are at the receiving end of the privacy and security support (Kropczynski et al., 2021a; Ghaiumy Anaraky et al., 2021), and are thus considered “caregivees”. However, community-based collaborative approaches may not take into account these unequal levels of expertise, as they have been primarily designed to support community members equally, with individuals giving and receiving feedback. Recently, Akter et al. (Akter et al., 2023b) conducted a field study using a community oversight app that allowed individuals in a community to help each other equally and found that users, in general, benefited from using the app. However, their work did not examine the presence or impact of potentially unequal roles within communities.
@@
This paper aims to address this gap, by investigating: 1) how caregivees vs. caregivers differ, 2) how mechanisms designed for collaborative community-based privacy and security support may affect caregivers and caregivees differently, and 3) if there are distinct differences in privacy behaviors between these two groups. To do this, we developed and deployed a community oversight app, called ”CO-oPS” that allows people of a trusted community to review one another’s mobile privacy and security decisions and provide advice and feedback on them. Our primary goal is to investigate how differently caregivers, who provide oversight, and caregivees, who receive the oversight, would be impacted by using the app. By providing insights into the varying effects of the CO-oPS app on these two unequal roles, we expect to gain generalizable knowledge that helps researchers and designers better understand the phenomena of privacy and security caregiving, how privacy-based interventions may effect these groups differently, and how to redesign future privacy-based interventions tailored to the unique needs of caregivers and caregivees. As such, we seek to answer the following high-level research questions:
@@ 2. Background
While these crowdsource-based solutions help individuals by raising their awareness about other users’ mobile privacy practices, they give little consideration to the trustworthiness of the information. Individuals might benefit more if the information comes from expert users who have expertise in mobile app privacy and security. Hence, Rashidi et al. (Rashidi et al., 2018) incorporated this idea of ”expert user” into their permission control framework titled ”DroidNet” which provided users recommendations based on the decisions made by the majority of expert users in the network. However, a large body of research work also demonstrated that individuals learn and change their digital privacy and security behaviors when they become aware of their close trusted circle’s privacy and security practices. For example, Das et al.’s studies (Das et al., 2014b, 2015) demonstrated the effectiveness of ”social proof” - i.e. being able to view how many friends in their social network use a specific security feature - and showed that individuals are more influenced to adopt a privacy and security feature when they witness adoption by others. Redmiles et al. also found that people trust and adopt privacy advice when it comes from their trusted close circles, e.g., family members, friends, and co-workers (Redmiles et al., 2016). In recent work, Mendel and Toch (Mendel and Toch, 2023) conducted a field experiment to evaluate a social support mechanism and found that users relied significantly more on the advice of close connections to learn about mobile security than community volunteers. Identifying the importance of social support in digital privacy and security, networked privacy researchers have investigated collaborative approaches. For instance, Wan et al. (Wan et al., 2020) designed a mobile app, titled AppMoD, to help individuals delegate their mobile privacy and security questions to an advisor - someone in their trusted social network. The advisor had the ability to make app permission decisions on behalf of the user. Chouhan et al. (Chouhan et al., 2019) proposed a community oversight framework to allow individuals in a trusted community, e.g., friends, family members, and co-workers, to help each other make their mobile privacy and security decisions together as a team.
@@
Communities can be heterogeneous when it comes to knowledge, behavior, skills, and roles in mobile privacy and security. Previous work examined how individuals in a community can be different in privacy and security knowledge, as well as take different roles in providing support. For example, Akter et al.’s studies (Akter et al., 2022, 2023a, 2024) revealed that although parents tended to take the role of supervising their families’ mobile online safety, privacy, and security, tech-savvy teens showed more concerns about the permissions granted on their parents’ phones, while parents were more focused on the kinds of apps their teens used. Researchers also examined how different age groups in communities vary in their decision-making about digital privacy and security. For example, Huang and Bashir (Huang and Bashir, 2018) identified that older adults, e.g., 65 years or older, tend to take fewer privacy protections than their younger peers, e.g., 55 to 64 years. However, in a more recent study, Anaraky et al. (Ghaiumy Anaraky et al., 2021) compared privacy decision making of young adults and older adults and found that young adults were more likely to disclose information based on their trust in the apps, while older adults, being careful, considered the pros and cons of each information disclosure. This body of research demonstrates that in communities, there may be some individuals who are more knowledgeable in making mobile privacy decisions and/or providing support to others (e.g., caregivers), and some might lack the knowledge and seek support from others (e.g., caregivees). Kropczynski et al. previously explored this concept of caregiving in the context of providing and receiving technological support within close, trusted circles, such as friends and family (Kropczynski et al., 2021b). Our study extends this concept to examine caregiving in the management of mobile privacy and security.
@@
The core idea behind collaborative community-based approaches (Wan et al., 2020; Chouhan et al., 2019; Aljallad et al., 2019) is to empower individuals within trusted communities to practice social support (DiGioia and Dourish, 2005; Das, 2016), enabling them to collaboratively assist one another in making both individual and collective privacy and security decisions. However, as suggested by Akter et al. (Akter et al., 2023b), these mechanisms are most effective when there are disparities in knowledge and expertise within the community, allowing those in need of support with mobile privacy and security decisions to benefit from the guidance of more knowledgeable community members. Despite this, the collaborative community-based mechanisms in the existing literature have not adequately considered the imbalance of expertise, caregiving roles, and the varying needs and perspectives (Badillo-Urquiola et al., 2020; Park et al., 2023) that accompany them. Therefore, the objective of our study was to investigate how these approaches, which treat all community members as equals, may or may not benefit caregivees who require additional support, as well as how they help caregivers support others. To examine these, we developed a collaborative community-based mobile privacy and security management app, CO-oPS, and conducted a study involving 22 groups of caregivers and caregivees who used the app within their groups. In the following section, we present a design overview of the CO-oPS app.
@@ 3. Design of the CO-oPS App
We developed the Android CO-oPS app based on the model of community oversight for privacy and security proposed by Chouhan et al. in (Chouhan et al., 2019). This app allows individuals in a trusted community, e.g., friends, families, and co-workers, to review one another’s apps installed and permissions granted, and provide or seek advice from one another, providing all community members the same interface and functionalities. Our app includes three main categories of features: 1) Co-monitoring, 2) Communication, and 3) Privacy.
@@ 4. Methods
The overall goal of our study is to examine the key differences in privacy and security behaviors between the individuals who provide oversight and those who receive when using a community oversight app. To achieve these goals, we recruited 22 small self-organized groups of people (2 to 6 individuals), who knew each other. In total, 101 participants were recruited across these groups, starting with primary contacts who then invited others within their social circles to join. Participants installed and used the CO-oPS Android app for four weeks with their group to exchange support and guidance on their privacy decisions. Participants also completed pre- and post-study surveys and were invited to participate in optional follow-up interviews. The study received approval from the Institutional Review Boards (IRBs) of the universities involved.
@@
Community Trust: In both the pre- and post-study surveys, we measured our participants’ perception of community trust. In this study, we created this construct based on the community oversight model framework that Chouhan et al. proposed in (Chouhan et al., 2019). We define community trust as an individual’s perception of trusting their community to keep their personal information, e.g., apps installed, private and to care for them to provide mobile privacy and security advice and guidance. The scale items are provided in Appendix A. Since our caregivees already received mobile privacy and security support from caregivers of their community before participating in the study, we expect caregivees will have more trust on their communities with their app usage information and will rely on others to receive advice and feedback on their mobile privacy and security. Moreover, as both caregivers and caregivees used the CO-oPS app to help each other with their mobile privacy and security decisions, we expect their community trust to increase at the end of the study. Therefore, we hypothesize a main effect of caregiving status prior to using the CO-oPS app, a main effect of the CO-oPS app intervention from pre- to post-study, and an interaction effect between the caregiving status and the CO-oPS app intervention:
@@
Community Collective Efficacy was defined by Carroll et al. (Carroll et al., 2005) as the capacity of a group or community for collaboratively performing a shared task. In our study, we utilized the pre-validated version of this construct which was used in Kropczynski et al.’s work (Kropczynski et al., 2021b). The community collective efficacy scale assessed how individuals perceived their community’s capacity to work together as a team on mobile privacy and security. Prior studies in the space of technology support (Kropczynski et al., 2021b) showed evidence that caregivers’ community collective efficacy was not different than the caregivees. Our study anticipates that the use of the CO-oPS app by caregivers and caregivees to exchange support and guidance on mobile privacy and security will enhance both groups’ collective efficacy in managing these issues within their community. Additionally, we expect that our caregivees would receive more help and guidance from the caregivers in making their mobile privacy and security decisions through using the CO-oPS app. Therefore, caregivees’ community collective efficacy will become stronger after using the CO-oPS app than the caregivers.
@@
Self-Efficacy was originally defined by Bandura (Bandura, 1982) as an individual’s perceived capacity to perform a task. In our study, we utilized the pre-validated version of this construct, which was used in Kropczynski et al.’s work (Kropczynski et al., 2021b). Their self-efficacy scale measured an individual’s perceived ability to manage their own privacy and security. Their work showed evidence that technology caregivers have higher self-efficacy than caregivees. In our study, since caregivers provided privacy and security support to their caregivees prior to participating in the study, we expect that our caregivers will also have higher self-efficacy for mobile privacy and security than the caregivees. Furthermore, since our caregivees used the CO-oPS app to learn from their community members’ mobile privacy and security behavior and received feedback and guidance from others on their own mobile privacy practices, we expect that our caregivees’ self-efficacy would become stronger after using the CO-oPS app than the caregivers. Therefore, we hypothesized that:
@@ 6. Discussion
Interestingly, we did not find significant differences between our caregivers and caregivees based on their demographic characteristics. Although this may partially be due to low sample-sizes across some groups (e.g., teens and older adults), anecdotally, we also found that older adults and youth in our study acted as caregivers. These findings run counter to common narratives in research (c.f.,  (Franz et al., 2019; Czaja et al., 2008; Kropczynski et al., 2021b; Akter et al., 2023a; Kiesler et al., 2000; Correa et al., 2015)), where older adults tend to seek support and younger generations provide the support to their families. Therefore, we give networked privacy and security researchers a word-of-caution – not to unintentionally stereotype caregivers/caregivees based on age or other demographics.
On the other hand, we hypothesized and confirmed that caregiving status differed significantly by power usage (i.e., higher for caregivers). This finding is important in that it serves to validate that individuals’ self-reported perceptions of who provides oversight aligned well with the level of power usage of these individuals, suggesting that caregiving for mobile privacy and security is related to the competence and desire to use technology to its fullest (Namara et al., 2021). This aligns with previous work (Akter et al., 2023b) which found that users often referred to tech savvy members of their community because they were more knowledgeable about mobile privacy and security. Therefore, future research could utilize power usage as a proxy measure to differentiate between caregivers and caregivees for mobile privacy and security within community-based settings – or it could simply ask individuals from whom they receive oversight as we did.
@@
In terms of the behavioral differences between our caregivers and caregivees, we only found statistically significant differences in the number of apps they hid from their community members. However, this is likely because we asked both our caregivers and caregivees to perform the same weekly tasks. However, qualitatively, we did see some differences in their privacy behaviors. Although our interview results showed a distinction between caregivers and caregivees in terms of initiating discussions to provide guidance and support and asking questions or seeking advice, our log results showed that both groups provided advice to the other caregivers and caregivees of their communities. This is because the collaborative community-based mechanisms of the CO-oPS app were designed to provide an equal footing for everyone in the community, regardless of their roles and expertise. Yet, our caregivers tended to review more apps, while caregivees changed relatively more app permissions based on the oversight received. In retrospect, caregivees providing advice may cause more harm than benefit, as they may lack the necessary skills to provide mobile privacy and security advice to others. This may also potentially cause tensions within the community, as caregivers might ignore their advice (Akter et al., 2023a) and the other caregivees may receive potentially wrong advice (Akter et al., 2022).
@@
Our results demonstrate how caregivers and caregivees are uniquely different in communities and how their perceptions differed when they used a community-based mobile privacy and security mechanism to exchange caregiving. The insights gained from the use of the CO-oPS app offer real-world applications not only for similar tools but also for the design of broader strategies aimed at helping users safeguard their privacy across various technological domains. These insights extend beyond mobile apps to other technology environments where users share personal information with third parties, such as smart home devices (Alghamdi et al., 2023, 2022), social media platforms (Bhagavatula et al., 2022; Ulusoy and Yolum, 2021) and websites (Mcdonald and Mentis, 2021). For example, collaborative community-based mechanisms can be adapted to assist individuals in making informed decisions about the data collection by IoT devices, as demonstrated in the work of Emami-Naeini et al. (Emami Naeini et al., 2018). Similarly, McDonald et al. (McDonald and Mentis, 2021) explored the design of technologies that enable collaboration among loved ones in managing cybersecurity, reflecting the broader applicability of community-driven approaches in enhancing privacy and security practices across different technological landscapes. Below we discuss the features and mechanisms that need to be in a collaborative oversight tool to support these differences of roles and expertise in communities.
@@
Addressing when Caregivers are Absent: In some cases, communities might lack technological expertise needed to provide care. To mitigate this problem, researchers and designers may need to explore ways to infuse external expertise in the community, as suggested by Chouhan et al. in (Chouhan et al., 2019). Another possibility is that rather than bringing external experts into a community, it might be advantageous to educate community members. In real-world settings, some caregivees may eventually learn to take the lead to provide oversight within their communities. Identifying these transitions over time is also important. Therefore, we recommend incorporating smart caregiving detection mechanisms into the application to recognize caregiving behaviors, such as sending messages containing privacy and security advice. This may require the integration of real-time messaging APIs to discern the intent behind messages, such as whether they are intended to seek or offer advice.
@@
We recognize several limitations of our study that should be addressed in future work. First, our sample was skewed toward younger adults with at least bachelor’s degrees. Our sample was also biased towards racial minorities with 72% of our participants being from Asian (primarily South Asian) descent. Further, our caregivers and caregivees were also generally equally distributed across these ethnicities, age groups, and education levels. Moreover, the co-monitoring feature of CO-oPS may have introduced conflicts among caregivees, as indicated by our RQ3 results, which showed that caregivees hid significantly more apps than caregivers. This issue could be further exacerbated in communities with power hierarchies, such as parent-child relationships, or among ethnic groups with lower privacy boundaries. Therefore, our results may not be generalizeable to populations of different ethnicities, age groups and education levels. Future work should explore caregiving with a community oversight mechanism among communities of more varied demographics and socio-economic status (Madden, 2017). One of the key limitations of our study was that we asked all of our participants to perform same set of tasks with CO-oPS, which resulted in non-significant differences in the behaviors of caregivees and caregivers. Additionally, the weekly tasks might have prompted caregivers and caregivees to adjust their app permissions or share oversight within their communities. In real-world scenarios, users might lack motivation for community oversight without incentives. Thus, we encourage further research into effective nudging mechanisms to incentivize users for community oversight and to take privacy protective actions for their own benefit.

### v2

@@ 1. Introduction
A recent Pew Research study (Vogels and Anderson, 2019) reported that most adults in the U.S. have significant knowledge gaps regarding their digital privacy and security. The report suggests a lack of understanding of how third party entities access personal information online (Auxier et al., 2019), making users susceptible to privacy and security breaches. This begs the question of how we might close this knowledge gap? Previous studies have explored solutions involving one party overseeing another, such as organizations using Mobile Device Management systems to secure employee data (Hayes et al., 2020) and parents monitoring teens through mobile online safety apps (Wisniewski et al., 2017). These studies suggested that collaborative approaches, rather than unilateral controls, can offer better privacy protection by fostering collective learning and discussion  (Akter et al., 2022b). Crowdsourcing has also been proposed as a method to improve individual’s mobile privacy decisions such as managing app permissions to safeguard personal information (Lin et al., 2012; Ismail et al., 2017; Zhang and Xu, 2016; Liu et al., 2019), though it often lacks reliability (Rashidi et al., 2018). An approach that has shown promise is to leverage the social influence of informal networks to infuse expertise and exert influence on privacy and security decisions (Mendel and Toch, 2017). For example, several studies have shown that people tend to trust and follow privacy advice from their trusted circles, often turning to friends and family for guidance on digital privacy and security topics (Mendel and Toch, 2017; Das et al., 2014a, b).
@@
Therefore, networked privacy researchers have called for community based approaches to co-manage digital privacy and security, where individuals in a trusted community can help one another equally make individual and collective privacy and security decisions (Wan et al., 2020; Chouhan et al., 2019; Aljallad et al., 2019). While this body of research confirms that people often rely on their social connections to inform their digital privacy and security decisions, it has yet to explore how these dynamics shift when relationships involve knowledge imbalances or power hierarchies, such as caring responsibilities for children or older adults. Our prior studies (Akter et al., 2022b, 2023b; Chouhan et al., 2019; Kropczynski et al., 2021b) have shown how communities often consist of hierarchical (i.e. uneven) levels of expertise and knowledge. Individuals with more expertise often take the role of “caregivers” to provide support and guidance to others who are less tech-savvy (Van Parys, 2019; Correa et al., 2015; Kiesler et al., 2000; Poole et al., 2009). Other individuals in the community are at the receiving end of the privacy and security support (Kropczynski et al., 2021a; Ghaiumy Anaraky et al., 2021), and are thus considered “caregivees”. However, community-based collaborative approaches may not take into account these unequal levels of expertise, as they have been primarily designed to support community members equally, with individuals giving and receiving feedback. Recently, we (Akter et al., 2023b) conducted a field study using a community oversight app that allowed individuals in a community to help each other equally and found that users, in general, benefited from using the app. Inspired by these positive results, we are now motivated to further investigate the impact of unequal roles within these collaborative environments.
@@
This paper aims to address this gap, by investigating: 1) how caregivees vs. caregivers differ, 2) how mechanisms designed for collaborative community-based privacy and security support may affect caregivers and caregivees differently, and 3) if there are distinct differences in privacy behaviors between these two groups. We developed and deployed a community oversight app, called ”CO-oPS” (Akter et al., 2022a) that allows people of a trusted community to review one another’s mobile privacy and security decisions and provide advice and feedback on them. Our primary goal in this paper is to investigate how differently caregivers, who provide oversight, and caregivees, who receive the oversight, would be impacted by using the app. By providing insights into the varying effects of the CO-oPS app on these two unequal roles, we expect to gain generalizable knowledge that helps researchers and designers better understand the phenomena of privacy and security caregiving, how privacy-based interventions may effect these groups differently, and how to redesign future privacy-based interventions tailored to the unique needs of caregivers and caregivees. As such, we seek to answer the following high-level research questions:
@@ 2. Background
While these crowdsource-based solutions help individuals by raising their awareness about other users’ mobile privacy practices, they give little consideration to the trustworthiness of the information. Individuals might benefit more if the information comes from expert users who have expertise in mobile app privacy and security. Hence, Rashidi et al. (Rashidi et al., 2018) incorporated this idea of ”expert user” into their permission control framework titled ”DroidNet” which provided users recommendations based on the decisions made by the majority of expert users in the network. However, a large body of research work also demonstrated that individuals learn and change their digital privacy and security behaviors when they become aware of their close trusted circle’s privacy and security practices. For example, Das et al.’s studies (Das et al., 2014b, 2015) demonstrated the effectiveness of ”social proof” - i.e. being able to view how many friends in their social network use a specific security feature - and showed that individuals are more influenced to adopt a privacy and security feature when they witness adoption by others. Redmiles et al. also found that people trust and adopt privacy advice when it comes from their trusted close circles, e.g., family members, friends, and co-workers (Redmiles et al., 2016). In recent work, Mendel and Toch (Mendel and Toch, 2023) conducted a field experiment to evaluate a social support mechanism and found that users relied significantly more on the advice of close connections to learn about mobile security than community volunteers. Identifying the importance of social support in digital privacy and security, networked privacy researchers have investigated collaborative approaches. For instance, Wan et al. (Wan et al., 2020) designed a mobile app, titled AppMoD, to help individuals delegate their mobile privacy and security questions to an advisor - someone in their trusted social network. The advisor had the ability to make app permission decisions on behalf of the user. In our prior work (Chouhan et al., 2019), we proposed a community oversight framework to allow individuals in a trusted community, e.g., friends, family members, and co-workers, to help each other make their mobile privacy and security decisions together as a team.
@@
Communities can be heterogeneous when it comes to knowledge, behavior, skills, and roles in mobile privacy and security. Previous work examined how individuals in a community can be different in privacy and security knowledge, as well as take different roles in providing support. For example, our prior studies (Akter et al., 2022b, 2023a, 2024) revealed that although parents tended to take the role of supervising their families’ mobile online safety, privacy, and security, tech-savvy teens showed more concerns about the permissions granted on their parents’ phones, while parents were more focused on the kinds of apps their teens used. Researchers also examined how different age groups in communities vary in their decision-making about digital privacy and security. For example, Huang and Bashir (Huang and Bashir, 2018) identified that older adults, e.g., 65 years or older, tend to take fewer privacy protections than their younger peers, e.g., 55 to 64 years. However, in a more recent study, Anaraky et al. (Ghaiumy Anaraky et al., 2021) compared privacy decision making of young adults and older adults and found that young adults were more likely to disclose information based on their trust in the apps, while older adults, being careful, considered the pros and cons of each information disclosure. This body of research demonstrates that in communities, there may be some individuals who are more knowledgeable in making mobile privacy decisions and/or providing support to others (e.g., caregivers), and some might lack the knowledge and seek support from others (e.g., caregivees). We have previously explored this concept of caregiving in the context of providing and receiving technological support within close, trusted circles, such as friends and family (Kropczynski et al., 2021b). This study extends this concept to examine caregiving in the management of mobile privacy and security.
@@
The core idea behind collaborative community-based approaches (Wan et al., 2020; Chouhan et al., 2019; Aljallad et al., 2019) is to empower individuals within trusted communities to practice social support (DiGioia and Dourish, 2005; Das, 2016), enabling them to collaboratively assist one another in making both individual and collective privacy and security decisions. However, one of our prior work (Akter et al., 2023b) suggested that these mechanisms are most effective when there are disparities in knowledge and expertise within the community, allowing those in need of support with mobile privacy and security decisions to benefit from the guidance of more knowledgeable community members. Despite this, the collaborative community-based mechanisms in the existing literature have not adequately considered the imbalance of expertise, caregiving roles, and the varying needs and perspectives (Badillo-Urquiola et al., 2020; Park et al., 2023) that accompany them. Therefore, the objective of this paper is to investigate how these approaches, which treat all community members as equals, may or may not benefit caregivees who require additional support, as well as how they help caregivers support others. To examine these, we developed a collaborative community-based mobile privacy and security management app, CO-oPS, and conducted a study involving 22 groups of caregivers and caregivees who used the app within their groups. In the following section, we present a design overview of the CO-oPS app.
@@ 3. Design of the CO-oPS App
We developed the Android CO-oPS app (Akter et al., 2022a) based on our proposed model of community oversight for privacy and security proposed by Chouhan et al. (Chouhan et al., 2019). This app allows individuals in a trusted community, e.g., friends, families, and co-workers, to review one another’s apps installed and permissions granted, and provide or seek advice from one another, providing all community members the same interface and functionalities. Our app includes three main categories of features: 1) Co-monitoring, 2) Communication, and 3) Privacy.
@@ 4. Methods
The overall goal of this paper is to examine the key differences in privacy and security behaviors between the individuals who provide oversight and those who receive when using a community oversight app. To achieve these goals, we recruited 22 small self-organized groups of people (2 to 6 individuals), who knew each other. In total, 101 participants were recruited across these groups, starting with primary contacts who then invited others within their social circles to join. Participants installed and used the CO-oPS Android app for four weeks with their group to exchange support and guidance on their privacy decisions. Participants also completed pre- and post-study surveys and were invited to participate in optional follow-up interviews. The study received approval from the Institutional Review Boards (IRBs) of the universities involved.
@@
Community Trust: In both the pre- and post-study surveys, we measured our participants’ perception of community trust. In this study, we created this construct based on our community oversight model framework that we proposed in (Chouhan et al., 2019). We define community trust as an individual’s perception of trusting their community to keep their personal information, e.g., apps installed, private and to care for them to provide mobile privacy and security advice and guidance. The scale items are provided in Appendix A. Since our caregivees already received mobile privacy and security support from caregivers of their community before participating in the study, we expect caregivees will have more trust on their communities with their app usage information and will rely on others to receive advice and feedback on their mobile privacy and security. Moreover, as both caregivers and caregivees used the CO-oPS app to help each other with their mobile privacy and security decisions, we expect their community trust to increase at the end of the study. Therefore, we hypothesize a main effect of caregiving status prior to using the CO-oPS app, a main effect of the CO-oPS app intervention from pre- to post-study, and an interaction effect between the caregiving status and the CO-oPS app intervention:
@@
Community Collective Efficacy was defined by Carroll et al. (Carroll et al., 2005) as the capacity of a group or community for collaboratively performing a shared task. In this study, we used a version of this construct that we validated in previous research (Kropczynski et al., 2021b). The community collective efficacy scale assessed how individuals perceived their community’s capacity to work together as a team on mobile privacy and security. Prior studies in the space of technology support (Kropczynski et al., 2021b) showed evidence that caregivers’ community collective efficacy was not different than the caregivees. Our study anticipates that the use of the CO-oPS app by caregivers and caregivees to exchange support and guidance on mobile privacy and security will enhance both groups’ collective efficacy in managing these issues within their community. Additionally, we expect that our caregivees would receive more help and guidance from the caregivers in making their mobile privacy and security decisions through using the CO-oPS app. Therefore, caregivees’ community collective efficacy will become stronger after using the CO-oPS app than the caregivers.
@@
Self-Efficacy was originally defined by Bandura (Bandura, 1982) as an individual’s perceived capacity to perform a task. In this study, we employed a pre-validated version of this construct from our prior work (Kropczynski et al., 2021b). Their self-efficacy scale measured an individual’s perceived ability to manage their own privacy and security. Their work showed evidence that technology caregivers have higher self-efficacy than caregivees. In our study, since caregivers provided privacy and security support to their caregivees prior to participating in the study, we expect that our caregivers will also have higher self-efficacy for mobile privacy and security than the caregivees. Furthermore, since our caregivees used the CO-oPS app to learn from their community members’ mobile privacy and security behavior and received feedback and guidance from others on their own mobile privacy practices, we expect that our caregivees’ self-efficacy would become stronger after using the CO-oPS app than the caregivers. Therefore, we hypothesized that:
@@ 6. Discussion
Interestingly, we did not find significant differences between our caregivers and caregivees based on their demographic characteristics. Although this may partially be due to low sample-sizes across some groups (e.g., teens and older adults), anecdotally, we also found that older adults and youth in our study acted as caregivers. These findings run counter to common narratives in research (c.f.,  (Franz et al., 2019; Czaja et al., 2008; Kropczynski et al., 2021b; Tanni et al., 2024; Akter et al., 2023a; Kiesler et al., 2000; Correa et al., 2015)), where older adults tend to seek support and younger generations provide the support to their families. Therefore, we give networked privacy and security researchers a word-of-caution – not to unintentionally stereotype caregivers/caregivees based on age or other demographics.
On the other hand, we hypothesized and confirmed that caregiving status differed significantly by power usage (i.e., higher for caregivers). This finding is important in that it serves to validate that individuals’ self-reported perceptions of who provides oversight aligned well with the level of power usage of these individuals, suggesting that caregiving for mobile privacy and security is related to the competence and desire to use technology to its fullest (Namara et al., 2021). As we have previously noted in (Akter et al., 2023b), CO-oPS users often referred to tech savvy members of their community because they were more knowledgeable about mobile privacy and security. Therefore, future research could utilize power usage as a proxy measure to differentiate between caregivers and caregivees for mobile privacy and security within community-based settings – or it could simply ask individuals from whom they receive oversight as we did.
@@
In terms of the behavioral differences between our caregivers and caregivees, we only found statistically significant differences in the number of apps they hid from their community members. Perhaps, this is likely because we asked both our caregivers and caregivees to perform the same weekly tasks. However, qualitatively, we did see some differences in their privacy behaviors. Although our interview results showed a distinction between caregivers and caregivees in terms of initiating discussions to provide guidance and support and asking questions or seeking advice, our log results showed that both groups provided advice to the other caregivers and caregivees of their communities. This is because the collaborative community-based mechanisms of the CO-oPS app were designed to provide an equal footing for everyone in the community, regardless of their roles and expertise. Yet, our caregivers tended to review more apps, while caregivees changed relatively more app permissions based on the oversight received. In retrospect, caregivees providing advice may cause more harm than benefit, as they may lack the necessary skills to provide mobile privacy and security advice to others. This may also potentially cause tensions within the community, as caregivers might ignore their advice (Akter et al., 2023a) and the other caregivees may receive potentially wrong advice (Akter et al., 2022b).
@@
Our results demonstrate how caregivers and caregivees are uniquely different in communities and how their perceptions differed when they used a community-based mobile privacy and security mechanism to exchange caregiving. The insights gained from the use of the CO-oPS app offer real-world applications not only for similar tools but also for the design of broader strategies aimed at helping users safeguard their privacy across various technological domains. These insights extend beyond mobile apps to other technology environments where users share personal information with third parties, such as smart home devices (Alghamdi et al., 2023, 2022a), social media platforms (Bhagavatula et al., 2022; Alghamdi et al., 2022b; Ulusoy and Yolum, 2021) and websites (Mcdonald and Mentis, 2021). For example, collaborative community-based mechanisms can be adapted to assist individuals in making informed decisions about the data collection by IoT devices, as demonstrated in the work of Emami-Naeini et al. (Emami Naeini et al., 2018). Similarly, McDonald et al. (McDonald and Mentis, 2021) explored the design of technologies that enable collaboration among loved ones in managing cybersecurity, reflecting the broader applicability of community-driven approaches in enhancing privacy and security practices across different technological landscapes. Below we discuss the features and mechanisms that need to be in a collaborative oversight tool to support these differences of roles and expertise in communities.
@@
Addressing when Caregivers are Absent: In some cases, communities might lack technological expertise needed to provide care. To mitigate this problem, researchers and designers may need to explore ways to infuse external expertise in the community, as we have previously suggested in (Chouhan et al., 2019). Another possibility is that rather than bringing external experts into a community, it might be advantageous to educate community members. In real-world settings, some caregivees may eventually learn to take the lead to provide oversight within their communities. Identifying these transitions over time is also important. Therefore, we recommend incorporating smart caregiving detection mechanisms into the application to recognize caregiving behaviors, such as sending messages containing privacy and security advice. This may require the integration of real-time messaging APIs to discern the intent behind messages, such as whether they are intended to seek or offer advice.
@@
We recognize several limitations of our study that should be addressed in future work. First, our sample was skewed toward younger adults with at least bachelor’s degrees. Our sample was also biased towards racial minorities with 72% of our participants being of Asian (primarily South Asian) descent. Further, our caregivers and caregivees were also generally equally distributed across these ethnicities, age groups, and education levels. Moreover, the co-monitoring feature of CO-oPS may have introduced conflicts among caregivees, as indicated by our RQ3 results, which showed that caregivees hid significantly more apps than caregivers. This issue could be further exacerbated in communities with power hierarchies, such as parent-child relationships, or among ethnic groups with lower privacy boundaries. Therefore, our results may not be generalizeable to populations of different ethnicities, age groups and education levels. Future work should explore caregiving with a community oversight mechanism among communities of more varied demographics and socio-economic status (Madden, 2017). One of the key limitations of our study was that we asked all of our participants to perform same set of tasks with CO-oPS, which resulted in non-significant differences in the behaviors of caregivees and caregivers. Additionally, the weekly tasks might have prompted caregivers and caregivees to adjust their app permissions or share oversight within their communities. In real-world scenarios, users might lack motivation for community oversight without incentives. Thus, we encourage further research into effective nudging mechanisms to incentivize users for community oversight and to take privacy protective actions for their own benefit.